Perkins Lighting wishes to employ a kanban in their new floor lamp production system. They will create production orders in the size of the kanban. For the floor lamp base, they have provided the following information: Daily demand 500 units Holding cost $15/unit/year Order cost $100/order Lead time 3 days Safety stock 500 units Find the size of the kanban and the number of kanbans required. Perkins believes in minimizing the total of the their ordering and holding costs.
Answer #1

Answer :

Annual Demand(A) = 500*365 = 182500

Carrying Cost(C) = $15

Ordering Cost(O) = $100

EOQ =  \sqrt{}2AO/C =  \sqrt{}(2*182500*100)15 = 1560 Units

Size of the order = 1560+500 = 2060 Units

Number of orders = 182500/2060 = 89 orders